# The Virtual Reality Koinos Method: Analyzing Virtual Reality Collaboration from the perspective of communication models

### v1

@@ 4.1.2 Procedure and participants
The experiment was presented to both participants together, explaining them that they would have to compare their impression between the ”human” and AI-controlled avatars. In the VR application, participants began by filling up a 2D grid assessing their emotion of Russel et al. [12], and completed a training of the task. The experimenter connected participants together, and they conducted the task. They ended the VR session by filling up a questionnaire that contained: a Fast Motion Sickness question of Keshavarz & Hecht [13] and the social presence questionnaire of Harms & Biocca [14] on a touch screen tablet. This study has recorded 52 participants, from different activity sector and different jobs. The average age was $M=32.44,SD=16.12$ and included 14 women and 38 men. This study was approved by an ethics committee.
@@ 4.1.3 Results
[…]
Figure 5: Co-presence results between Human and AI avatar experimental conditions with 52 participants divided equally between the two conditions.
[…]
@@ 4.1.4 Discussion
The study was analyzed regarding the AI social representation of the online survey, and shows that co-presence and behavior were influenced by it. Indeed, the social representation of AI seems relevant to explain the difference of co-presence level and the absence of difference in task performance. ”AI” is seen as a help, as an assistant, participants may have expected an efficient behavior, and saw it as non-human entity, resulting in a lower co-presence level than in the Human condition. In addition, correlation tests did reveal three factors that can be related to communication (e.g., Perceived Behavior Interdependence, Looking at the avatar’s body, Perceived Message Understanding).
@@ 4.2 VR experiment 2: Does the avatar appearance really matter?
Given the results of the first experiment, varying the social representations of the virtual collaborators with a congruent avatar appearance, we explored another aspect of social interactions: the impact of the avatar appearance on social interactions. Using the same task as in the first experiment, the main condition was the change of avatar appearance with the same virtual partner; participants completed the task together and knowing that they were collaborating together.
@@ 4.2.1 Experimental design
A similar experimental design was used as the one for the previous VR experiment. Four avatar appearances were compared: one meshed avatar, the most used according to Weidner, et al. [15], and three point-cloud avatars. Point-cloud avatars were used in the objective to induce lower social bias associated to the avatar appearance, based on (Hudson & Hurter [16] and Yee & Bailenson [17], while allowing an anthropomorphic representation and reproduction of human behavior. Due to organization constraints, participants tested two avatars, divided into two groups. One group tested in random order, the meshed (M, Figure 6 up-left) avatar and the avatar with the highest number of point avatar (HP, 70 000 points, Figure 6 up-right), allowing to recognize the shape of the meshed avatar. The second group had tested in random order, the medium number of point (M, 10 000 points, Figure 6 down-left) and the lowest number of point (LP, 60 points, Figure 6 down-right) avatars. Before the beginning of the task, participants experienced an embodiment phase in front of a virtual mirror, following five instructions of 40 seconds each.
@@ 4.2.2 Procedure and participants
A similar procedure was conducted as in experiment 1. However, some differences can be noted: participants were told collaborating together in both VR sessions, an embodiment phase was added in the VR application, as well as an embodiment questionnaire (Gonzalez-Franco & Peck, [18]). This study has recorded 48 participants, mainly students from business school, engineering school and computer science programs. The average age was $M=21.17,SD=3.34$ and included 16 women and 32 men. This study was approved by an ethics committee.
@@ 4.2.3 Results
[…]
@@
[…]
@@ 4.2.4 Discussion
This study was analyzed with communication models, using the ABX model of Newcomb [3] and the Cooperative Model of Human Communication of Tomasello [4]. The participants having a strong common ground, the communication cues were limited and the co-presence level was high for each avatar. A ceiling effect was observed on the co-presence level results. Therefore, we suggest that the avatar appearance did not impact the social interaction, thus had less impact than the common ground. This analysis led us to suggest to investigate the impact of communication cues on the social interaction perceived quality, as well as investigate abstract avatars that could reproduce human behavior with lower social bias than meshed avatars.
@@ 4.3 Implementing the Koinos Method on the VR experiments
The application of the communication models on the second experiment allowed us to better understand the social interaction dynamics in this collaborative task, and also helped us propose that the avatar appearance may have less impact in the social interaction than a strong common ground.
We proposed to apply the same analysis on the first VR experiment, that uses the same VR application. The participants communicated through their avatars and the VR hardware, Virtual Environment and User Interface. However, contrary to the second experiment, participants did not know with who they completed the task. The experimenter explained that they would complete the task either with a human, a student in another lab that they did not know, or with an avatar controlled by an AI. In real, participants completed the task together. The results indicated lower co-presence levels in the AI condition than with the Human. We first apply the common ground evaluation grid for the Human condition (Table 2):
Only one possibility on the two choices in each part:
@@
With the human partner, due to the collaborative task, we propose that the common ground is based on an immediate perceptual environment, on a top-down process and on explicit common cultural knowledge, the human partner being presented as a student in another laboratory of the same country.
With the AI partner, the use of the grid may be not relevant. Indeed, the AI entity is a non-human entity. Therefore, it is possible to say that AI is seen as an entity that it is possible to cooperate with to complete a task (related to the social presence presented in the study of the first experiment), but we cannot argue if this entity is perceived as equal as a human to cooperate. Indeed, with a human, the Theory of Minds of Gallagher & Frith [2] explains that human beings are considered as entities having their own thoughts, beliefs, values, goals, are therefore relevant as cooperative entities and correspond to the Cooperative Model of Human Communication of Tomasello [4]. With the actual knowledge and literature, we may say that the common ground is null with a non-human entity. However, this assumption requires further study, considering that humans are increasingly accustomed to interacting with non-human entities, such as AI, and that their behavior is becoming more similar to human behavior. Regarding the results of AI social representations online survey and the resulting data of the experiment, we can suppose that social representations of the AI entity impact the interaction. However, social representations are certainly linked to common ground: for instance, in a given situation, even if the common ground can be strong, the social representations of the other person or entity influence behaviors. In other words, the common ground can be considered null with the AI entity and higher with the human entity, whereas the communication cues are the same in both cases and the same in the second experiment, therefore limited with only animation of the hands and the head movements.
From the two experiments, the co-presence level is higher in the second experiment than in the human condition of the first. Despite differential elements in the experimental procedures, we can suggest that the common ground is stronger in the second experiment than in the first; participants knew with whom they would complete the task, while they thought that they did not know their partner in the first experiment. We apply the common ground evaluation grid on the second experiment (Table 3):
Only one possibility on the two choices in each part:
@@
A score of 6 results of the use of the common ground evaluation grid for the second experiment, while a score of 4 for the first experiment in Human condition and a (supposed) score of 0 for the AI condition. The co-presence level is then correlated to these scores, despite the use of the same limited communication cues. Therefore, we can suppose that with a high common ground between participants, it is not necessary to have complex communication cues for participants to experience a high co-presence level. On the other hand, having a lower common ground is associated to a lower co-presence level, and even lower with a non-human partner.
@@
In other words, when participants perceived the interdependence of behavior, they also experienced a high co-presence level, and similar for PMU. Based on communication theories, the behavioral interdependence and message understanding are essential to experience smooth social interaction. Therefore, we can suppose that this task proposed relevant communication cues (exclusively non-verbal) regarding the objectives of the task, and that it engaged participants together.
@@ 5 General Discussion: Toward a Co-presence equation?
The application of the proposed Koinos Method allows us to identify a pattern that could link co-presence, common ground and communication cues, and are summarized in the Table 4.
@@ 6 Limited
$Co-Presence=Common\ ground\ +\ Communication\ cues$ (1)
The two experiments allowed us to verify the variation of common ground, however the communication cues were the same. The results of the second experiment suggest that one of the experimental conditions used by Gamelin [19] and Yu [20] has no effect (appearance), implying that the observed differences in their studies may be attributed to the second condition (the tracking system). Our second experiment varied the avatar appearance (meshed and point-cloud avatars) with the same tracking system, and a high co-presence value resulted for each avatar. Although an experiment is needed to validate this theory, we can suggest that the tracking system have causes the difference in co-presence in both study of Gamelin [19] and Yu [20]. If it is confirmed, our proposed the equation of co-presence can be validated: with the same common ground, changing the communication cues can influence the co-presence level. Regarding non-verbal communication, we can suppose that a more precise, detailed, human-like tracking and animation could increase the co-presence value, and therefore the social interaction perceived experience. Studies about co-presence with agents point in this direction (Sterna, et al. [21] and Strojny et al. [22]).
The analyzes lead us to the main result of this study: with the same communication cues, a stronger common ground results in a higher co-presence level. Therefore, it is possible to use the co-presence equation to increase the social interaction perceived, depending on the task, on the relation and the nature of the entity of interactors, and the communication cues. It is possible to use the common ground evaluation grid to evaluate it at first sight, and adapt the necessary communication cues, along with the interactions needed to complete the task or the objective of the VR application.
@@ 5.1 Limitations and Future works
The establishment of the Co-Presence equation is mainly based on the already known impact of communication cues on co-presence, and on the two VR experiments presented in this paper, therefore to validate it, a deeper analysis has to be done on different task type and different social interaction situations. Further studies could be done on the common ground evaluation grid as well as describe communication cues elements.
From our analyses of the VR experiments and the literature, future work could be on three different aspects, exploring different types of abstract avatars, studying the impact of the tracking system on social interaction and performed advanced studies of social representation of agents.
@@ 5.2 Usage for developers
This study aimed at using communication models to better understand social interaction in VR, especially regarding the co-presence level, that reflects the perceived quality of the social interaction.
The result of the analysis of the VR experiment led us to propose an equation that can allow to predict the co-presence level, and identify what elements can be changed in the VR application to enable higher co-presence levels. We can propose to developers to use the common ground evaluation grid to first evaluate how strong the common ground between users can be; and in a second step evaluate if a complex communication is necessary (if the common ground is low). Finally, we can suggest that the design of the avatar and the VR application has to be relevant to both complete the task or the objective, and allow to communicate with others. The second study seems to indicate that the avatar appearance is not primary on the common ground, if the latter is strong.
@@ 6 Conclusion
The main result of this study indicates that using communication models to better understand co-presence in VR seems relevant. We propose a methodology that adapts communication models to the specificities of VR systems, with the aim of analyzing the dynamics of social interactions occurring in VR: the Koinos Method. The analysis of the two VR experiments using the Koinos Method indicates that the avatar appearance does not seems prior on co-presence level and task completion. In addition, our results show that the prejudices and ideas about the virtual collaborator guide the behavior and the perceived experience of social interaction.
Secondly, the analysis of the VR experiments with communication models leads us to suggest that the co-presence level could depend on the common ground between participants and the communication cues. Co-presence being considered as an indication of the social interaction perceived quality, our results show that with a strong common ground, complex communication cues are not necessary. Conversely, with a light or null common ground, with a non-human partner such as an agent, communication cues have to be more complex in order to experience higher co-presence. In this case, studying the social representation associated to the non-human partner could help identify which factors could bring this non-human entity more ”human”. In the case of a social interaction with a human entity, an evaluation of the strength of the common ground seems relevant to determine if complex communication cues have to be included; in the objective to bring the perceived quality of social interactions closer to the FtF interactions. In this case, we proposed a co-presence equation along with the Common Ground Evaluation Grid, allowing to predict the co-presence level and to identify which factors could help upgrade this level and the social interaction quality.
These results highlight the impact of the relationship between participants on co-presence. Previous studies on so-presence usually vary technological elements without precising the relationship between participants, especially about the common ground shared. Therefore, we suggest that researchers consider using the Common Ground Evaluation Grid presented in this study to assess the level of common ground between participants in future studies.
@@
Research was conducted as part of a PhD thesis fully funded by a French regional ground.

### v2

@@ 4.1.2 Procedure and participants
The experiment was presented to both participants together, explaining that they would have to compare their impressions of the Human-controlled avatar and AI-controlled avatar conditions. In the VR application, participants began by filling in a 2D grid assessing their emotion of Russel et al. [russell1989affect], and completed a training of the task. The experimenter connected participants together, and they conducted the task. They ended the VR session by filling out a questionnaire that contained: a Fast Motion Sickness question of Keshavarz & Hecht [keshavarz2011validating] and the social presence questionnaire of Harms & Biocca [harms2004internal] on a touch screen tablet. This study has recorded 52 participants from various sectors and jobs roles. The average age was $M=32.44,SD=16.12$ and included 14 women and 38 men. This study was approved by an ethics committee from Bourgogne Franche-Comté University, n°CERBFC-2023-05-25-028.
@@ 4.1.3 Results
The data were first tested for normality and homoscedasticity, which revealed that the use of non-parametric tests was relevant. The first main result of this study is the difference in co-presence level between the Human ($M=1.87,SD=1.21$) and AI avatars experimental conditions ($M=1.41,SD=1.26$, with $Z=2.21,p=.027$) from a Wilcoxon signed-rank test (Figure 5).
Figure 5: Co-presence results between Human and AI avatars experimental conditions with 52 participants divided equally between the two conditions.
The second result is the absence of differences in the task performance, either in score and time completion, between the AI ($M=191.34,SD=50.14$) and Human ($M=191.96,SD=50.04$) avatars conditions, $Z=.30,p=.76$. Finally, Spearman correlation tests indicate that the three variables showing the highest Correlation Factor ($CF$) with co-presence are Perceived Behavioral Interdependence ($CF=.545,p<.001$) (Harms & Biocca, [harms2004internal]), Looking at the Avatar’s Body ($CF=.486,p<.001$), and Perceived Message Understanding ($CF=.442,p<.001$) (Harms & Biocca, [harms2004internal]).
@@ 4.1.4 Discussion
The study was analyzed regarding the AI social representation of the online survey, and it shows that co-presence and behavior were influenced by it. The co-presence level is significantly lower in the AI-controlled avatar condition than with the Human-controlled avatar. The social representation of AI seems relevant to explain this result. An AI is seen as a help, as an assistant, participants may have expected an efficient behavior to complete the task, and saw it as a non-human entity. Therefore, the feeling of being with someone in VR with an AI-controlled avatar is significantly lower than with a Human-controlled avatar, even told that this human is not known.
In addition, correlation tests did reveal three factors that can be related to communication (e.g., Perceived Behavior Interdependence, Looking at the avatar’s body, Perceived Message Understanding).
@@ 4.2 VR experiment 2: Does the avatar appearance really matter?
Given the results of the first experiment, in which we varied the social representations of the virtual collaborators by using a congruent avatar appearance, we explored another aspect of social interactions: the impact of the avatar appearance on social interactions. Using the same task as in the first experiment, the main condition involved changing the avatar appearance with the same virtual partner; participants completed the task together, knowing that they were collaborating, in contrast to the first experiment.
@@ 4.2.1 Experimental design
A similar experimental design was used as the one for the previous VR experiment. Four avatar appearances were compared: one meshed avatar, the most used according to Weidner et al. [weidner2023systematic], and three point-cloud avatars. Point-cloud avatars were used in the objective to induce lower social bias associated with the avatar’s appearance, based on (Hudson & Hurter [hudson2016avatar] and Yee & Bailenson [yee2007proteus], while allowing an anthropomorphic representation and reproduction of human behavior. Due to organizational constraints, participants tested two avatars, divided into two groups. One group tested in random order, the meshed (M, Figure 6 up-left) avatar and the avatar with the highest number of point avatar (HP, 70 000 points, Figure 6 up-right), allowing to recognize the shape of the meshed avatar. The second group had tested in random order, the medium number of points (M, 10 000 points, Figure 6 down-left) and the lowest number of points (LP, 60 points, Figure 6 down-right) avatars. Before the task began, participants underwent an embodiment phase in front of a virtual mirror, following five instructions of 40 seconds each.
@@ 4.2.2 Procedure and participants
A similar procedure was conducted as in experiment 1. However, some differences can be noted: participants were told to collaborate together in both VR sessions, an embodiment phase was added in the VR application, as well as an embodiment questionnaire (Gonzalez-Franco & Peck, [gonzalez2018avatar]). This study has recorded 48 participants, mainly students from business school, engineering school, and computer science programs. The average age was $M=21.17,SD=3.34$ and included 16 women and 32 men. This study was approved by an ethics committee from Bourgogne Franche-Comté University, n°CERBFC-2024-06-03-022.
@@ 4.2.3 Results
The data were first tested for normality and homoscedasticity, which revealed that the use of non-parametric tests was relevant. First, a high co-presence level was noted for each avatar, with a ceiling effect (Figure 7). The average values and Standard deviation of each avatar are: $M=2.33,SD=.94$ for the M avatar, $M=2.33,SD=.90$ for the HP avatar, $M=2.25,SD=1.18$ for the MP avatar and $M=2.22,SD=1.27$ for the LP avatar. A mixed linear model analysis did not show any effect of the avatar condition on the co-presence level (F=.03,p=.9F=.03,p=.97).
@@
Secondly, the results did not represent significant difference in task performance regarding the avatar condition, either during the training ($F=.50,p=.68$) or during the task ($F=.30,p=.82$).
A Spearman correlation test did not reveal any significant correlation the between embodiment sub-categories (e.g., Body Ownership, Agency and Motor Control, and External appearance) results and the co-presence level. However, the observed ceiling effect may have impacted these results. It is interesting to note that a significant effect is present between the four avatars regarding the embodiment sub-categories in “Agency and Motor control” ($F=3.68,p=.02$) and “External appearance” ($F=7.60,p<.001$).
Finally, the Spearman test indicates that the three variables showing the highest Correlation Factor ($CF$) with co-presence are Perceived Behavioral Interdependence ($CF=.641,p<.001$), Emotional valence before VR ($CF=.398,p<.001$) and Perceived Message Understanding ($CF=.347,p<.005$).
@@ 4.2.4 Discussion
This study was analyzed with communication models, using the ABX model of Newcomb [newcomb1953approach] and the Cooperative Model of Human Communication of Tomasello [tom2008]. The participants having a strong common ground, the communication cues were limited, and the co-presence level was high for each avatar. A ceiling effect was observed on the co-presence level results. Therefore, the avatar’s appearance did not impact the social interaction, and thus had less impact than the common ground. This study compared abstract avatars, with a point cloud representation, with human body representation, using the same tracking system and avatar animation (inverse kinematic). This analysis suggests investigating the impact of communication cues on the social interaction perceived quality, as well as investigating abstract avatars that could reproduce human behavior with lower social bias than meshed avatars.
@@ 4.3 Implementing the Koinos Method on the VR experiments
The application of the communication models in the second experiment allowed us to better understand the social interaction dynamics in this collaborative task. Also, it helped us propose that the avatar’s appearance may have less impact on the social interaction than a strong common ground.
We propose applying the same analysis to the first VR experiment, which utilizes the same VR application. The participants communicated through their avatars and the VR hardware, Virtual Environment, and User Interface. However, contrary to the second experiment, participants did not know with whom they completed the task. The experimenter explained that they would complete the task either with a human, a student in another lab that they did not know, or with an avatar controlled by an AI. In reality, participants completed the task together. The results indicated lower co-presence levels in the AI-controlled avatar condition than in the Human-controlled avatar condition. We first apply the common ground evaluation grid for the Human condition (Table 2):
Only one possibility in each of the three columns:
@@
With the human partner, due to the collaborative task, we propose that common ground is based on an immediate perceptual environment, on a top-down process, and on explicit common cultural knowledge, the human partner being presented as a student in another laboratory of the same country.
With the AI partner, the use of the grid may not be relevant. Indeed, the AI entity is a non-human entity. Therefore, AI is seen as an entity with which it is possible to cooperate to complete a task (related to the social presence presented in the first experiment). However, we cannot argue whether this entity is perceived as equal to a human in terms of cooperation. Indeed, with a human, the Theory of Minds of Gallagher & Frith [gallagher2003functional] explains that human beings are considered as entities having their own thoughts, beliefs, values, goals, are therefore relevant as cooperative entities and correspond to the Cooperative Model of Human Communication of Tomasello [tom2008]. With the actual knowledge and literature, we may say that the common ground is null with a non-human entity. However, this assumption requires further study, considering that humans are increasingly accustomed to interacting with non-human entities, such as AI, and that their behavior is becoming more similar to human behavior.
Regarding the results of the AI social representations online survey and the resulting data of the experiment, social representations of the AI entity impact the interaction. However, social representations are indeed linked to common ground: for instance, in a given situation, even if the common ground can be strong, the social representations of the other person or entity influence behaviors. In other words, the common ground can be considered null with the AI entity and higher with the human entity, whereas the communication cues are the same in both cases and the same in the second experiment, therefore limited to only animation of the hands and the head movements.
From the two experiments, the co-presence level is higher in the second experiment than in the first human condition. Despite differential elements in the experimental procedures, the common ground is stronger in the second experiment than in the first; participants knew with whom they would complete the task, while they thought that they did not know their partner in the first experiment. We apply the common ground evaluation grid on the second experiment (Table 3):
Only one possibility in each of the three columns:
@@
A score of 6 results from the use of the common ground evaluation grid for the second experiment, while a score of 4 for the first experiment in the Human-controlled avatar condition, and a (supposed) score of 0 for the AI-controlled avatar condition. The co-presence level is then correlated to these scores, despite the use of the same limited communication cues. Therefore, with a high common ground between participants, it is not necessary to have complex communication cues for participants to experience a high co-presence level. On the other hand, having a lower common ground is associated with a lower co-presence level, and even lower with a non-human partner.
@@
In other words, when participants perceived the interdependence of behavior, they also experienced a high level of co-presence, and this was similar for PMU. Based on communication theories, behavioral interdependence and message understanding are essential to experiencing smooth social interaction. Therefore, this task proposed relevant communication cues (exclusively non-verbal) regarding the objectives of the task, and it engaged participants together.
@@ 5 General Discussion: Toward a Co-presence equation?
The application of the proposed Koinos Method allows us to identify a pattern that seems to link co-presence, common ground and communication cues, and are summarized in the Table 4.
@@ 6 Limited
CoPresence=Commonground+CommunicationcuesCo-Presence=Common\ ground\ +\ Communication\ cues (1)
The two experiments allowed us to verify the variation of common ground; however, the communication cues were the same. The results of the second experiment suggest that one of the experimental conditions used by Gamelin [gamelin2021point] and Yu [yu2021avatars] has no effect (appearance), implying that the observed differences in their studies may be attributed to the second condition (the tracking system). Our second experiment varied the avatar appearance (meshed and point-cloud avatars) with the same tracking system, and a high co-presence value resulted for each avatar. Although an experiment is needed to validate this theory, we can suggest that the tracking system has caused the difference in co-presence in both studies of Gamelin [gamelin2021point] and Yu [yu2021avatars]. If it is confirmed, our proposed equation of co-presence can be validated: with the same common ground, changing the communication cues can influence the co-presence level. Regarding non-verbal communication, a more precise, detailed, human-like tracking and animation could increase the co-presence value, and therefore the social interaction perceived experience. Studies about co-presence with agents point in this direction (Sterna et al. [sterna2024behavioral] and Strojny et al. [strojny2020moderators]).
The analyses lead us to the main result of this study: with the same communication cues, a stronger common ground results in a higher co-presence level. Therefore, although this theory needs to be validated, it is possible to use the co-presence equation to increase the social interaction perceived, depending on the task, the relation, the nature of the entity of interactors, and the communication cues. It is possible to use the common ground evaluation grid to evaluate it at first sight, and adapt the necessary communication cues, along with the interactions needed to complete the task or the objective of the VR application.
@@ 5.1 Limitations and Future works
The establishment of the Co-Presence equation is mainly based on the already known impact of communication cues on co-presence, and on the two VR experiments presented in this paper; therefore, to validate it, a deeper analysis has to be done on different task types and different social interaction situations. Further studies could be done on the common ground evaluation grid, as well as describe communication cues elements.
From our analyses of the VR experiments and the literature, future work could be on three different aspects: exploring different types of abstract avatars, studying the impact of the tracking system on social interaction, and performing advanced studies of social representation of agents.
@@ 5.2 Usage for developers
This study aimed to use communication models to better understand social interaction in VR, especially regarding the co-presence level, which reflects the perceived quality of the social interaction.
The result of the analysis of the VR experiment led us to propose an equation that can allow us to predict the co-presence level and identify what elements can be changed in the VR application to enable higher co-presence levels. We can propose to developers to use the common ground evaluation grid first to evaluate how strong the common ground between users can be, and in a second step, evaluate if complex communication is necessary (if the common ground is low). Finally, the design of the avatar and the VR application has to be relevant to both complete the task or the objective, and allow communication with others. The second study indicates that the avatar’s appearance is not primary on the common ground, if the latter is strong.
@@ 6 Conclusion
The main result of this study indicates that using communication models to better understand co-presence in VR seems relevant. We propose a methodology that adapts communication models to the specificities of VR systems, with the aim of analyzing the dynamics of social interactions occurring in VR: the Koinos Method. The analysis of the two VR experiments using the Koinos Method indicates that the avatar’s appearance does not affect the co-presence level and task completion in this particular task. In addition, our results show that the prejudices and ideas about the virtual collaborator guide the behavior and the perceived experience of social interaction.
Secondly, the analysis of the VR experiments with communication models suggests that the co-presence level depends on the common ground between participants and the communication cues. Co-presence being considered as an indication of the social interaction perceived quality, our results show that with a strong common ground, complex communication cues are not necessary. Conversely, with a light or null common ground, with a non-human partner such as an agent, communication cues have to be more complex in order to experience higher co-presence. In this case, studying the social representation associated with the non-human partner could help identify which factors could bring this non-human entity more ”human”. In the case of a social interaction with a human entity, an evaluation of the strength of the common ground seems relevant to determine if complex communication cues have to be included, with the objective of bringing the perceived quality of social interactions closer to the FtF interactions. In this case, we proposed a co-presence equation along with the Common Ground Evaluation Grid, allowing us to predict the co-presence level and to identify which factors could help upgrade this level and the social interaction quality.
These results highlight the impact of the relationship between participants on co-presence. Previous studies on co-presence usually vary technological elements without specifying the relationship between participants, especially regarding the common ground shared. Therefore, we suggest that researchers consider using the Common Ground Evaluation Grid presented in this study to assess the level of common ground between participants in future studies.
@@
This work was supported by the French region of Bourgogne-Franche-Comté under the researcher-entrepreneur program (ICE) grant 2022 Y_23815.